# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 2 $\{0\}$$\{1\}$$\{2\}$
Conjunction: Suppose that $v_{\Delta}(\alpha\land\beta)=2$. Then, $\Box(\alpha\land\beta)\in\Delta$ and so $\Box\alpha\land\Box\beta\in\Delta$, since $\Box(\alpha\land\beta)\to(\Box\alpha\land\Box\beta)$ is a theorem in $\mathcal{H_{S4}}$. This implies that $\Box\alpha\in\Delta$ and $\Box\beta\in\Delta$, hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$. Thus, $v_{\Delta}(\alpha\land\beta)=2\in\{2\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Now, assume that $v_{\Delta}(\alpha\land\beta)=1$. Then, $\Box(\alpha\land\beta)\notin\Delta$ but $\alpha\land\beta\in\Delta$ and so $\alpha\in\Delta$ and $\beta\in\Delta$ but either $\Box\alpha\notin\Delta$ or $\Box\beta\notin\Delta$ (since $(\Box\alpha\land\Box\beta)\to\Box(\alpha\land\beta)$ is a theorem in $\mathcal{H_{S4}}$). That is, $v_{\Delta}(\alpha),v_{\Delta}(\beta)\in\{1,2\}$ and either $v_{\Delta}(\alpha)=1$ or $v_{\Delta}(\beta)=1$. Then, $v_{\Delta}(\alpha\land\beta)=1\in\{1\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\land\beta)=0$ then $\alpha\land\beta\notin\Delta$, hence $\alpha\notin\Delta$ or $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=0$ or $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\land\beta)=0\in\{0\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
Implication: Suppose that $v_{\Delta}(\alpha\to\beta)=2$. Then, $\Box(\alpha\to\beta)\in\Delta$ and so $\alpha\to\beta\in\Delta$, by axiom T. This implies that either $\alpha\notin\Delta$ or $\beta\in\Delta$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$. Thus, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. If $\beta\in\Delta$, suppose first that $\Box\alpha\in\Delta$. Then, by axiom K it follows that $\Box\beta\in\Delta$. Hence, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$ and so $v_{\Delta}(\alpha\to\beta)=2\in\{2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\Box\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\neq 2$ and $v_{\Delta}(\beta)\neq 0$. From this, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\to\beta)=1$. Then, $\Box(\alpha\to\beta)\notin\Delta$ but $\alpha\to\beta\in\Delta$ and so $\alpha\notin\Delta$ or $\beta\in\Delta$. In addition, $\Box\beta\notin\Delta$, since $\Box\beta\to\Box(\alpha\to\beta)$ is a theorem in $\mathcal{H_{S4}}$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$ and $v_{\Delta}(\beta)\in\{0,1\}$. Thus, $v_{\Delta}(\alpha\to\beta)=1\in\{1,2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\beta\in\Delta$ then $v_{\Delta}(\beta)=1$ and so $v_{\Delta}(\alpha\to\beta)=1\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\to\beta)=0$ then $\alpha\to\beta\notin\Delta$, hence $\alpha\in\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)\in\{1,2\}$ and $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\to\beta)=0\in\{0\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
It is analogous to the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Lemma 3.9 and Theorem 3.10], but now by using our Lemma 3.12 instead of [4, Lemma 3.7]. ∎
Now, Definitions 3.5 and 3.6 can be easily adapted to $\mathcal{M}^{\prime}_{S4}$, obtaining so the set $PLV^{\prime}$ of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over finite sets of formulas. However, it will be convenient to give a bit more detailed presentation of these sets, by specifying explicitly the domain of each valuation.
@@
[…]
@@
[…]
@@
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $v(\beta)=2$. If $\Lambda\in FCS(\Sigma)$ then
$\begin{array}[]{lll}PLV(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV(\Lambda)$}\big{)}\}.\end{array}$
As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $PV(\Lambda)$ is finite, hence the definition above is not cyclic.
The next step is to show analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}^{\prime}_{S4}$. In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$.
@@
Consider the predicate $\mathsf{P}_{\Lambda}(v,w,\alpha)$ introduced in Remark 3.17. Clearly,
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}^{\prime}_{S4}$}\big{)}\}.\end{array}$
@@
It is an adaptation of the proof of [4, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin PLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=2\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$ and, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)=0$ implies that $w(\beta)\neq 2$ for some $\beta\in\Gamma$. Hence, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)\not\in\mathcal{D}$ implies that $w(\Box\beta)\notin\mathcal{D}$ for some $\beta\in\Gamma$. By contraposition, this means that $\{\Box\beta\mid\beta\in\Gamma\}\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\alpha$. By Theorem 3.13, $\{\Box\beta\mid\beta\in\Gamma\}\vdash_{S4}\alpha$. Observe that $\nvdash_{S4}\alpha$: otherwise, $v(\alpha)=2$, by Corollary  3.11.1. Hence, there exist $\beta_{1},\ldots,\beta_{n}\in\Gamma$ such that $\vdash_{S4}\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))$. By necessitation rule, $\vdash_{S4}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. By Theorem 3.11, $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. But $v(\Box\beta_{i})=2$ for $1\leq i\leq n$ and $v(\alpha)=1$, hence $v(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))=1$. This means that $v(\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))))=0$, a contradiction. From this we infer that $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. ∎
@@
[…]
@@
Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Sigma)$ such that $\textsf{c}_{1}(\alpha_{i})\leq\textsf{c}_{1}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}_{1}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Sigma)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
@@
It is an adaptation of the proof of [4, Lemma 3.14]. In this case, more modifications are required because of the conjunction and disjunction connectives, and by the fact of considering a fixed set $\Lambda$.
@@
Suppose first that $\Lambda=For(\Sigma)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for every $n$. By Definition 3.15 of partial valuation, $\tilde{v}_{p}\in Val(\mathcal{M}^{\prime}_{S4})={\mathcal{L}^{\prime}}_{0}^{S4}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for a given $n\geq 0$, and let $\alpha$ such that $\mathcal{L}^{\prime S4}_{n}(\alpha)\subseteq\mathcal{D}$. In particular, $\tilde{v}_{p}(\alpha)\in\mathcal{D}$. Suppose that $\tilde{v}_{p}(\alpha)=1$. By Definition 3.18 of partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}^{\prime}_{S4}$ such that $w(\alpha)=0$. But this is a contradiction, since $w\in\mathcal{L}^{\prime S4}_{n}$. Then $\tilde{v}_{p}(\alpha)=2$ and so $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}^{\prime}_{S4}$ and the result clearly holds.
@@
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Sigma)$. Consider an enumeration of $For(\Sigma)$ as in Remark 3.22. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. Observe that $\Lambda\cup\mathcal{P}\in CS(\Sigma)$ and, for every $n\in\omega$, $\Lambda_{n}\in CS(\Sigma)$ for $\Lambda_{n}:=\Lambda\cup For(\Sigma)_{n}$. In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}_{1}(\alpha_{m})=n+1$, then $\Lambda_{n}^{m}\in CS(\Sigma)$, where $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$. Clearly $\Lambda_{n+1}=\cup_{k\in\omega}\Lambda_{n}^{\omega\cdot(n+1)+k}$.
@@
Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=0$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.

### v4

@@ 2 $\{0\}$ $\{1\}$ $\{2\}$
Conjunction: Suppose that $v_{\Delta}(\alpha\land\beta)=2$. Then, $\Box(\alpha\land\beta)\in\Delta$ and so $\Box\alpha\land\Box\beta\in\Delta$, since $\Box(\alpha\land\beta)\to(\Box\alpha\land\Box\beta)$ is a theorem in $\mathcal{H}_{S4}$. This implies that $\Box\alpha\in\Delta$ and $\Box\beta\in\Delta$, hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$. Thus, $v_{\Delta}(\alpha\land\beta)=2\in\{2\}=\land^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Now, assume that $v_{\Delta}(\alpha\land\beta)=1$. Then, $\Box(\alpha\land\beta)\notin\Delta$ but $\alpha\land\beta\in\Delta$ and so $\alpha\in\Delta$ and $\beta\in\Delta$ but either $\Box\alpha\notin\Delta$ or $\Box\beta\notin\Delta$ (since $(\Box\alpha\land\Box\beta)\to\Box(\alpha\land\beta)$ is a theorem in $\mathcal{H}_{S4}$). That is, $v_{\Delta}(\alpha),v_{\Delta}(\beta)\in\{1,2\}$ and either $v_{\Delta}(\alpha)=1$ or $v_{\Delta}(\beta)=1$. Then, $v_{\Delta}(\alpha\land\beta)=1\in\{1\}=\land^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\land\beta)=0$ then $\alpha\land\beta\notin\Delta$, hence $\alpha\notin\Delta$ or $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=0$ or $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\land\beta)=0\in\{0\}=\land^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
Implication: Suppose that $v_{\Delta}(\alpha\to\beta)=2$. Then, $\Box(\alpha\to\beta)\in\Delta$ and so $\alpha\to\beta\in\Delta$, by axiom T. This implies that either $\alpha\notin\Delta$ or $\beta\in\Delta$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$. Thus, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. If $\beta\in\Delta$, suppose first that $\Box\alpha\in\Delta$. Then, by axiom K it follows that $\Box\beta\in\Delta$. Hence, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$ and so $v_{\Delta}(\alpha\to\beta)=2\in\{2\}={\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\Box\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\neq 2$ and $v_{\Delta}(\beta)\neq 0$. From this, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\to\beta)=1$. Then, $\Box(\alpha\to\beta)\notin\Delta$ but $\alpha\to\beta\in\Delta$ and so $\alpha\notin\Delta$ or $\beta\in\Delta$. In addition, $\Box\beta\notin\Delta$, since $\Box\beta\to\Box(\alpha\to\beta)$ is a theorem in $\mathcal{H}_{S4}$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$ and $v_{\Delta}(\beta)\in\{0,1\}$. Thus, $v_{\Delta}(\alpha\to\beta)=1\in\{1,2\}={\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\beta\in\Delta$ then $v_{\Delta}(\beta)=1$ and so $v_{\Delta}(\alpha\to\beta)=1\in{\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\to\beta)=0$ then $\alpha\to\beta\notin\Delta$, hence $\alpha\in\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)\in\{1,2\}$ and $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\to\beta)=0\in\{0\}={\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
It is analogous to the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [12, Lemma 3.9 and Theorem 3.10], but now by using our Lemma 6.4 instead of [12, Lemma 3.7]. ∎
Now, Definitions 5.4 and 5.5 can be easily adapted to $\mathcal{M}^{\prime}_{S4}$, obtaining so the set $PLV^{\prime}$ of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over finite sets of formulas. However, it will be convenient to give a bit more detailed presentation of these sets, by specifying explicitly the domain of each valuation.
@@
$CS(\Sigma)=\{\Lambda\subseteq For(\Sigma)\mid\mbox{ $\Lambda$ is non-empty and closed under subformulas}\},$
@@
  • if $#\in\{\neg,\Box\}$ and $\alpha\in\Lambda$ then $\tilde{v}_{p}(\#\alpha)\in\#^{S4}(\tilde{v}_{p}(\alpha))$;
  • if $#\in\{\to,\vee,\land\}$ and $\alpha\#\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\#\beta)\in\#^{\prime S4}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
@@
Given $v,w$ and $\alpha$ let $\mathsf{P}^{4}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $v(\beta)=2$. If $\Lambda\in FCS(\Sigma)$ then
$\begin{array}[]{lll}PLV(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV(\Lambda)$}\big)\}.\end{array}$
As observed in Remark 5.6 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $PV(\Lambda)$ is finite, hence the definition above is not cyclic. To be more precise, the following result can be proven:
Proposition 6.10 (Existence and uniqueness of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
Let $\Lambda\in FCS(\Sigma)$. Then, there exists a unique set $PLV(\Lambda)\subseteq PV(\Lambda)$ satisfying the condition of Remark 6.9. Moreover, $PLV(\Lambda)\neq\emptyset$.
Proof.
The proof can be easily adapted from the corresponding one for Proposition 9.7 for $\mathbf{IPL}$ to be given in Section 9. ∎
The next step is to show analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}^{\prime}_{S4}$. In order to do this it is considered, adapting [12, Definition 3.12], the intermediary notion of partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$.
@@
Consider the predicate $\mathsf{P}^{4}_{\Lambda}(v,w,\alpha)$ introduced in Remark 6.9. Clearly,
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}^{\prime}_{S4}$}\big)\}.\end{array}$
@@
It is an adaptation of the proof of [12, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin PLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=2\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$ and, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)=0$ implies that $w(\beta)\neq 2$ for some $\beta\in\Gamma$. Hence, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)\not\in\mathcal{D}$ implies that $w(\Box\beta)\notin\mathcal{D}$ for some $\beta\in\Gamma$. By contraposition, this means that $\{\Box\beta\mid\beta\in\Gamma\}\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\alpha$. By Theorem 6.5, $\{\Box\beta\mid\beta\in\Gamma\}\vdash_{S4}\alpha$. Observe that $\nvdash_{S4}\alpha$: otherwise, $v(\alpha)=2$, by Corollary 6.2.1. Hence, there exist $\beta_{1},\ldots,\beta_{n}\in\Gamma$ such that $\vdash_{S4}\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))$. By necessitation rule, $\vdash_{S4}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. By Theorem 6.2, $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. But $v(\Box\beta_{i})=2$ for $1\leq i\leq n$ and $v(\alpha)=1$, hence $v(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))=1$. This means that $v(\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))))=0$, a contradiction. From this we infer that $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. ∎
@@
The complexity $\textsf{c}_{1}(\alpha)$ of a formula $\alpha\in For(\Sigma)$ is defined as follows: $\textsf{c}_{1}(p)=0$ if $p$ is a propositional variable; $\textsf{c}_{1}(\neg\alpha)=\textsf{c}_{1}(\Box\alpha)=\textsf{c}_{1}(\alpha)+1$; and $\textsf{c}_{1}(\alpha\#\beta)=\textsf{c}_{1}(\alpha)+\textsf{c}_{1}(\beta)+1$, for $#\in\{\to,\vee,\land\}$. Let $For(\Sigma)_{n}=\{\alpha\in For(\Sigma)\mid\textsf{c}_{1}(\alpha)\leq n\}$ for every $n\geq 0$.
@@
Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Sigma)$ such that $\textsf{c}_{1}(\alpha_{i})\leq\textsf{c}_{1}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}_{1}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Sigma)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 0 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
Based on the enumeration of $For(\Sigma)$ given in the previous remark, consider the following sets:
Definition 6.16.
Let $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) be an enumeration of $For(\Sigma)$ as in Remark 6.15. For every $n,m\in\omega$ let ${\Lambda}_{n}:=\Lambda\cup For(\Sigma)_{n}$; ${\Lambda}_{n}^{0}:={\Lambda}_{n}$; and ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$.
Remark 6.17.
By definition, ${\Lambda}_{n}^{m+1}={\Lambda}_{n}\cup\{\alpha_{\omega\cdot(n+1)},\alpha_{\omega\cdot(n+1)+1},\ldots,\alpha_{\omega\cdot(n+1)+m}\}$, and $\alpha_{\omega\cdot(n+1)+m}\in\Lambda$ if and only if ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}$. Hence, ${\Lambda}_{n}^{m+1}$ is obtained from ${\Lambda}_{n}$ by adding the first $m+1$ formulas (of the given enumeration) with complexity $n+1$, and so ${\Lambda}_{n}^{m}$ adds at most $m$ formulas to ${\Lambda}_{n}$, for $m\in\omega$. It is worth noting that ${\Lambda}_{n}^{m}\in CS(\Sigma)$ for every $n,m\in\omega$; in particular, ${\Lambda}_{n}\in CS(\Sigma)$ for every $n\in\omega$. Clearly ${\Lambda}_{n}^{m}\subseteq{\Lambda}_{n}^{m+1}$, ${\Lambda}_{n+1}=\bigcup_{m\in\omega}{\Lambda}_{n}^{m}$, and $For(\Sigma)=\bigcup_{n\in\omega}{\Lambda}_{n}$.
@@
It is an adaptation of the proof of [12, Lemma 3.14]. In this case, more modifications are required because of the conjunction and disjunction connectives, and by the fact of considering a fixed set $\Lambda$.
@@
Suppose first that $\Lambda=For(\Sigma)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for every $n$. By Definition 6.7 of partial valuation, $\tilde{v}_{p}\in Val(\mathcal{M}^{\prime}_{S4})={\mathcal{L}^{\prime}}_{0}^{S4}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for a given $n\geq 0$, and let $\alpha$ such that $\mathcal{L}^{\prime S4}_{n}(\alpha)\subseteq\mathcal{D}$. In particular, $\tilde{v}_{p}(\alpha)\in\mathcal{D}$. Suppose that $\tilde{v}_{p}(\alpha)=1$. By Definition 6.11 of partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}^{\prime}_{S4}$ such that $w(\alpha)=0$. But this is a contradiction, since $w\in\mathcal{L}^{\prime S4}_{n}$. Then $\tilde{v}_{p}(\alpha)=2$ and so $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}^{\prime}_{S4}$ and the result clearly holds by taking $v:=\tilde{v}_{p}$.
@@
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Sigma)$. Consider an enumeration of $For(\Sigma)$ as in Remark 6.15. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. More precisely, with terminology as in Definition 6.16 and from the observations in Remark 6.17, an extension $v_{n}^{m}$ of $\tilde{v}_{p}$ (i.e., $\tilde{v}_{p}\subseteq v_{n}^{m}$) will be defined for every $(n,w)\in\omega\times\omega$ such that $v_{n}^{m}\in PLV^{\prime}({\Lambda}_{n}^{m})$ and $v_{i}^{j}\subseteq v_{n}^{m}$ if $(i,j)\leq(n,m)$, where: $(i,j)\leq(n,m)$ iff $i\leq n$ or $i=n$ and $j\leq m$ (and where $\omega\times\omega$ denotes, as usual, the Cartesian product of $\omega$ with itself).
@@
Base $n=m=0$: Observe that ${\Lambda}_{0}={\Lambda}_{0}^{0}=\Lambda\cup\mathcal{P}$. Define $v_{0}^{0}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v_{0}^{0}(p)=0$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $\tilde{v}_{p}\subseteq v_{0}^{0}$ and $v_{0}^{0}\in PLV^{\prime}({\Lambda}_{0})=PLV^{\prime}({\Lambda}_{0}^{0})$. Let $v_{0}:=v_{0}^{0}$.
Inductive step: Assume that $v_{i}^{j}(\alpha)$ was defined for every $\alpha\in{\Lambda}_{i}^{j}$ such that $v_{i}^{j}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, $v_{i}^{j}\in PLV^{\prime}({\Lambda}_{i}^{j})$, $v_{i}^{j}\subseteq v_{k}^{r}$ if $(i,j)\leq(k,r)\leq(n,m)$, for given $n,m\geq 0$ (Induction Hypothesis, IH). Now it will be shown how to extend $v_{n}^{m}$ to a function $v_{n}^{m+1}$ with domain ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$. That is, it will be shown how to define a value for $\alpha_{\omega\cdot(n+1)+m}$ while preserving the values assigned by $v_{n}^{m}$, in such a manner that $v_{n}^{m+1}\in PLV^{\prime}({\Lambda}_{n}^{m+1})$. To start with, define $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ for every $\beta\in{\Lambda}_{n}^{m}$.